OpenStudy (anonymous):

Question 1. What kind of lines are these, perpendicular or parallel? Question 2. What is the slope of the first line you drew? Question 3. What is the slope of the second line you drew? Question 4. What is the relationship between the slopes of two parallel lines? On the same graph, plot the points (-1, 5) and (5, 1) on the grid paper using one of your colored pencils and connect them. Question 5. What is the slope of the line you just drew? Question 6. What is its relationship to the first line you drew, perpendicular or parallel? Question 7. What is the relationship between the

OpenStudy (anonymous):

wow that is lot questions lol i will try my best

OpenStudy (anonymous):

Lol sorry

OpenStudy (anonymous):

I have to go run to the store real quick. can you keep going till I get back?

OpenStudy (anonymous):

yeah sure

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

All right lets get start! 1) These are parallel. 2) Slope of first line = 3/2 3) Slope of second line = 3/2 4) The slopes of the two lines are equal. the equation of the third line is y = (-2/3)x + 13/3 5) Slope of the third line = -2/3 6) Perpendicular 7) The product of the slopes of two lines must be -1 Hope i helped. Anyway i gotta go hope some people can answer your questions ;)